A particle moves around the x axis with constant acceleration of 4ms^-2.at Time t =0 it is at x =5m and has velocity v =3ms^-1.Find the position and velocity at Time t =2 secs.Where is the body when its velocity is 5ms^-1
1
Expert's answer
2019-01-23T10:28:06-0500

The position of the body

x(t)=x_0+v_0 t+(at^2)/2

x(2)=5+3×2+(4×2^2)/2=19 m

The velocity of the body

v(t)=v_0+at

v(2)=3+4×2=11 m/s

v(t)=3+4t=5

t=0.5 s